OpenStudy (anonymous):

Find the polynomial function f(x) of lowest degree which has zeros 2i (of multiplicity of 2) and 3 (of mult. of 1) and for which f(1)=5.

OpenStudy (freckles):

with real coefficients?

OpenStudy (anonymous):

whatever that means

OpenStudy (anonymous):

do you know a simple quadratic that has one zero of \(2i\)>

OpenStudy (freckles):

oh actually (x-2i)^2

OpenStudy (freckles):

lol

OpenStudy (freckles):

wait is that question to me @satellite73 or to the person

OpenStudy (freckles):

lol

OpenStudy (freckles):

but if he wants real coefficients which I'm assume he does he will need both (x-2i)^2 and (x+2i)^2 to be factors of the given poly

OpenStudy (freckles):

that is why the i'm asking about the real coefficients thing...

OpenStudy (anonymous):

Yes, using real coef. so that would make the lowest degree 5?

OpenStudy (anonymous):

of course they are supposed to be real the quadratic with zero of \(2i\) is \(x^2+4\) since if \[x=2i\\x^2=-4\\ x^2+4=0\]

OpenStudy (anonymous):

yes, degree 5

OpenStudy (freckles):

polys can have complex coefficients

OpenStudy (freckles):

but yeah if you want real coefficients you can say you have x=2i twice which means you have the factor x-2i twice what I mean is (x-2i)^2 will be a factor of your poly and since we want real coefficents you will neeed also x=-2i twice or the factor x+2i twice so you will have (x-2i)^2(x+2i)^2 which you can expand to see you will have all real coefficients then if x=3 you know what else is a factor of your poly? once you found that I will on and say how to use the f(1)=5 thing

OpenStudy (freckles):

or you can use what @satellite73 said above

OpenStudy (freckles):

x^2+4=0 has solutions x=2i or x=-2i so to get those zeros twice how many (x^2+4) 's do you need in your poly?

OpenStudy (anonymous):

I'm at (x-2i)^2(x+2i)^2(x-3) and setting f(1) would make it (1-2i)^2(1+2i)^2(-2)

OpenStudy (anonymous):

whoa nelly too much work !

OpenStudy (freckles):

one sec though ...so yeah you do have all of that (x-2i)^2(x+2i)^2(x-3) but you need one more thing since you have the condition f(1)=5 you need a constant in which to solve for so that when you replace x with 1 you get output 5 so you have \[f(x)=a(x-2i)^2(x+2i)^2(x-3)\]

OpenStudy (anonymous):

multiply out and deal with real numbers

OpenStudy (anonymous):

a would be 5

OpenStudy (anonymous):

no one wants to compute what you wrote \[f(x)=a(x^2+4)^2(x-3)\] then put \(f(1)=5\) and solve for \(a\)

OpenStudy (freckles):

\[f(1)=a(1-2i)^2(1+2i)^2(1-3) \\ 5=a[(1-2i)(1+2i)]^2(-2) \\ 5=a[1-4i^2]^2(-2) \\ 5=a[1+4]^2(-2)\]

OpenStudy (freckles):

so a isn't 5

OpenStudy (anonymous):

Ok, I see. So 5=-50a?

OpenStudy (freckles):

right and then divide both sides by -50

OpenStudy (freckles):

and reduce fraction

OpenStudy (freckles):

couple of things I want to mention if they weren't clear above I used law of exponents \[a^2 \cdot b^2 =(a \cdot b)^2\] and I also used that easy formula for multiplying conjugates \[(a+bi)(a-bi)=a^2-b^2 \cdot i^2=a^2+b^2\] or you might know it better as \[(a-b)(a+b)=a^2-b^2\]

OpenStudy (anonymous):

Yes, it makes sense where it came from.